Grade 12Magical Mathematics[Interesting Approach]

I am little bit confused in a question that If there are two Children in a family then probability that both are boys What Should be its Ans can anyone please explain

Sagar Rangari
10 Years agoGrade 12

2 Answers

Devesh Paliwal
10 Years ago
Here, Sample Space = {B-B, B-G, G-G} 
                      so, n(s)=3        
         Events having both boys= {B-B}
                      so, n(E)= 1
 
Thus,
                   PE = n(E)/n(S)
                   PE = 1/3                             Ans.
Devesh Paliwal
10 Years ago
I am Sorry above one is incorrect..
 
The right one is
 
Here, Sample Space = {B-B, B-G, G-B, G-G}
 
so, n(s)=4        
         Events having both boys= {B-B}
                      so, n(E)= 1
 
Thus,
                   PE = n(E)/n(S)
                   PE = 1/4                             Ans.
 
This one is correct